# The Limited Impact of Medical Adaptation of Large Language and Vision-Language Models

### v1

@@ C.2 Default LLM Prompt Formats
Here, we share the default prompt formats that we use for each LLM, using MMLU (Clinical Knowledge) (Hendrycks et al., 2021) as a running example. We denote the system prompt in red, any few-shot examples in green, and the question being asked of the model in purple.
For models that do not have a specific system prompt and prompt format designed for closed-ended medical QA (see Section C.1), we use a manually designed prompt format by default. This includes all of the general-domain LLMs. For example, in the 1-shot setting, the default prompt for non-instruction-tuned models is as follows:
@@
For instruction-tuned models, which typically expect a specific conversational format, we apply the above format to each “user” query and “assistant” response and remove the ‘###’ and ‘Answer:’ tags. For example, the input prompt to Llama-3-70B-Instruct is as follows:
@@
<|begin_of_text|>
<|start_header_id|> system <|end_header_id|>
@@
<|system|>: You are a helpful medical assistant created by M42 Health in the UAE.
<|prompter|>: Question: Glycolysis is the name given to the pathway involving the conversion of:
@@ C.3 Default VLM Prompt Formats
In this section, we share the default prompt formats that we use for each general-domain/medical VLM, using VQA-RAD (Lau et al., 2018) as a running example. We denote the system prompt in red, any few-shot examples in green, and the question being asked of the model in purple. By default, we show the format used in the 1-shot setting.
@@
Appendix D Additional Details on Supervised Fine-Tuning
In this section, we provide additional details on the supervised fine-tuning experiments detailed in Section 3.2. As discussed, we fine-tune all models except Med-Flamingo-9B and Open-Flamingo-9B with LoRA adapters, and mainly search over the LoRA ranks and learning rates. For Med-Flamingo-9B and Open-Flamingo-9B, we fine-tune the gated cross-attention layers, the Perceiver resampler (Jaegle et al., 2021), and the embeddings of the <image> and <|endofchunk|> tokens, and mainly search over the learning rates and weight decay coefficients. Below, we summarize all of the hyperparameter search spaces that we perform a grid search over for each model.
LoRA fine-tuning details for LLMs.
For all fine-tuning runs, we fix the $\alpha$ hyperparameter in LoRA to 16 and search over the ranks $r=[16,32,64]$ and the learning rates $\eta=[10^{-5},2\times 10^{-5},5\times 10^{-5},10^{-4},2\times 10^{-4}]$, resulting in 15 trials per model and dataset. Across the different models and datasets, we adjust the per-device batch size $B$ as follows to ensure that no memory errors are encountered during training:
  • 8B-parameter LLMs:
    $B=8$ (MedQA (4 Options), PubMedQA, CASI: Sense, MIMIC-III: Sense)
    $B=4$ (EHRNoteQA)
    $B=2$ (i2b2 datasets, when $r\in\{16,32\}$)
    $B=1$ (i2b2 datasets, when $r=64$)
    $B=16$ (Otherwise)
  • 7B-parameter LLMs:
    $B=16$ (PubMedQA, CASI Sense Diambiguation)
    $B=4$ (i2b2 datasets)
    $B=32$ (Otherwise)
For each run, we train the model for a maximum of 10 epochs and apply early stopping regularization with a patience of 1. We track the validation cross-entropy loss on the output tokens after every epoch, and save the checkpoint that achieved the lowest validation loss during each fine-tuning run. We then select the model that achieved the lowest validation loss across all trials to use for final evaluation on the test set. We train all models using Stage 2 of DeepSpeed ZeRO (Rajbhandari et al., 2020). On the MedMCQA and PubMedQA datasets, which are the largest textual QA datasets considered in our evaluations (see Table A1), we use 4 NVIDIA A6000 GPUs per trial to increase the effective batch size and speed up training. For all other textual QA datasets, we use 2 NVIDIA A6000 GPUs per trial.
LoRA fine-tuning details for LLaVA-Med-v0-7B and LLaVA-v0-7B.
For all fine-tuning runs, we fix the $\alpha$ hyperparameter in LoRA to 16 and search over the ranks $r=[16,32,64]$, the learning rates $\eta=[10^{-5},2\times 10^{-5},5\times 10^{-5}]$, and LoRA dropout $p=[0,0.1]$, resulting in 18 trials per model and dataset. For all datasets, we fix the per-device batch size $B$ to 32. For each run, we train the model for a maximum of 10 epochs and apply early stopping regularization with a patience of 1. We track the validation cross-entropy loss on the output tokens after every epoch, and save the checkpoint that achieved the lowest validation loss during each fine-tuning run. We then select the model that achieved the lowest validation loss across all trials to use for final evaluation on the test set. We train all models using Stage 2 of DeepSpeed ZeRO (Rajbhandari et al., 2020). For all visual QA datasets, we use 2 NVIDIA A6000 GPUs per trial.
Fine-tuning details for Med-Flamingo-9B and Open-Flamingo-9B.
For all fine-tuning runs, we mainly search over the learning rates $\eta=[10^{-5},2\times 10^{-5},5\times 10^{-5}]$ and the weight decay coefficients $\lambda=[0,0.05,0.1]$, resulting in 9 trials per model and dataset. For all datasets, we fix the per-device batch size $B$ to 16. For each run, we train the model for a maximum of 10 epochs and apply early stopping regularization with a patience of 1. We track the validation cross-entropy loss on the output tokens after every epoch, and save the checkpoint that achieved the lowest validation loss during each fine-tuning run. We then select the model that achieved the lowest validation loss across all trials to use for final evaluation on the test set. We train all models using PyTorch FSDP (Zhao et al., 2023), following the official implementation provided by Awadalla et al. (2023). For all visual QA datasets, we use 6 NVIDIA A6000 GPUs per trial.
@@ Appendix D Additional Details on Supervised Fine-Tuning
Appendix E Zero-/Few-Shot Prompting Evaluation Results with Constrained Decoding
Here, we provide the constrained decoding results for the zero-/few-shot prompting evaluations described in Section 3.1 (see “Models” in Section 3 for a description on constrained decoding). In Appendices E.1E.3, corresponding to Sections 4.1.14.1.3, wer show the results on the textual medical knowledge QA, textual clinical note QA, and visual medical QA datasets, when model-specific prompt selection is performed for each model independently. In Appendix E.4, corresponding to Section 4.2.3, we show the aggregate results on how the win, tie, and loss rates (%) of the medical models are affected by the inclusion (or lack thereof) of model-specific prompt selection and statistical testing.
E.1 Evaluation of Medical LLMs on Textual Medical Knowledge QA (Section 4.1.1)
Figure E1: After independently selecting the best prompt format and examples for each model, medical LLMs do not consistently show a statistically significant improvement over their general-domain counterparts on textual medical knowledge QA tasks, in both (a) zero-shot and (b) 3-shot settings. The top row of each panel shows the absolute exact-match accuracies on the test set, and the bottom row of each panel shows the relative exact-match accuracies along with 95% confidence intervals derived via bootstrapping on the test set (see Section 3). Here, we show the results when model predictions are generated via constrained decoding. The results for greedy decoding are shown in Figure 3.
@@ Appendix E Zero-/Few-Shot Prompting Evaluation Results with Constrained Decoding
Table E1: The zero-shot and 3-shot win/tie/loss rates (%) of all medical LLMs on textual medical knowledge QA, after independently optimizing the prompt for each model. For each medical model, we boldface the win rate if it wins more than it loses to its general-domain base model, and vice versa. Here, we show the results when model predictions are generated via constrained decoding. The results for greedy decoding are shown in Table 2.
@@
OpenBioLLM-70B (Pal and Sankarasubbu, 2024) 7.7 76.9 15.4 23.1 53.8 23.1
MediTron-70B (Chen et al., 2023) 30.8 46.2 23.1 15.4 69.2 15.4
Clinical-Camel-70B (Toma et al., 2023) 18.2 72.7 9.1 9.1 72.7 18.2
OpenBioLLM-8B (Pal and Sankarasubbu, 2024) 0 53.8 46.2 7.7 69.2 23.1
MediTron-7B (Chen et al., 2023) 23.1 76.9 0 7.7 92.3 0
BioMistral-7B (Labrak et al., 2024) 30.8 69.2 0 7.7 92.3 0
BioMedGPT-LM-7B (Luo et al., 2023) 7.7 84.6 7.7 7.7 69.2 23.1
Aggregate 16.9 68.6 14.5 11.2 74.1 14.7
In Figure E1, we show the absolute and relative exact-match accuracies achieved by the medical and general-domain LLMs on the textual medical knowledge QA datasets, from zero-shot (Figure E1(a)) and 3-shot prompting (Figure E1(b)), respectively. In Table E1, we also show the win, tie, and loss rates (%) of the medical LLMs, where win rate refers to the proportion of QA datasets where a medical model shows a statistically significant improvement over its base model. For each medical model in Table E1, we boldface the win rate if it wins more than it loses to its general-domain base model, and vice versa. We exclude the results for Clinical-Camel-70B on both versions of MedQA, and the Med42 models on all textual medical knowledge QA datasets, as they have already been trained on these datasets as part of medical DAPT (see Table 1).
Overall, we observe that our findings discussed in Section 4.1.1 also hold in the constrained decoding setting. In particular, we observe that after performing model-specific prompt selection for each model independently,
@@
    more recent general-domain models (e.g., Llama-3-8B) tend to show stronger performance even without any medical adaptation via DAPT;
@@
    the majority of medical LLMs show limited improvement over their base models, reaching a tie in most cases (aggregate zero-shot/3-shot tie rates of 68.6%/74.1%).
@@
E.2 Evaluation of Medical LLMs on Textual Medical Knowledge QA (Section 4.1.2)
Figure E2: After independently selecting the best prompt format and examples for each model, medical LLMs do not consistently show a statistically significant improvement over their general-domain counterparts on textual clinical note QA tasks, in both (a) zero-shot and (b) 3-shot settings. In the 3-shot setting, we exclude the results on the EHRNoteQA and i2b2 datasets due to the limited context window sizes of the medical LLMs that we evaluate (see Section 3.1). The top row of each panel shows the absolute exact-match accuracies on the test set, and the bottom row of each panel shows the relative exact-match accuracies along with 95% confidence intervals derived via bootstrapping on the test set (see Section 3). Here, we show the results when model predictions are generated via constrained decoding. The results for greedy decoding are shown in Figure 4.
Table E2: The zero-shot and 3-shot win, tie, and loss rates (%) of all medical LLMs on textual clinical note QA, after independently optimizing the prompt for each model. For each medical model, we boldface the win rate if it wins more than it loses to its general-domain base model, and vice versa. Here, we show the results when model predictions are generated via constrained decoding. The results for greedy decoding are shown in Table 3.
@@
Med42-v2-70B (Christophe et al., 2024b) 12.5 75.0 12.5 0 66.7 33.3
OpenBioLLM-70B (Pal and Sankarasubbu, 2024) 0 75.0 25.0 0 33.3 66.7
Med42-v1-70B (Christophe et al., 2024a) 87.5 0 0 66.7 33.3 0
MediTron-70B (Chen et al., 2023) 12.5 75.0 12.5 0 33.3 66.7
Clinical-Camel-70B (Toma et al., 2023) 87.5 12.5 0 33.3 0 66.7
Med42-v2-8B (Christophe et al., 2024b) 37.5 37.5 25.0 66.7 0 33.3
OpenBioLLM-8B (Pal and Sankarasubbu, 2024) 37.5 25.0 37.5 66.7 0 33.3
MediTron-7B (Chen et al., 2023) 0 66.7 33.3 33.3 66.7 0
BioMistral-7B (Labrak et al., 2024) 0 12.5 87.5 33.3 0 66.7
BioMedGPT-LM-7B (Luo et al., 2023) 50.0 12.5 37.5 0 33.3 66.7
Aggregate 32.5 39.1 28.3 33.3 23.3 43.3
In Figure E2, we show the absolute and relative exact-match accuracies achieved by the medical and general-domain LLMs on the textual medical knowledge QA datasets, from zero-shot (Figure E2(a)) and 3-shot prompting (Figure E2(b)), respectively. In Table E2, we also show the win, tie, and loss rates (%) of the medical LLMs, where win rate refers to the proportion of QA datasets where a medical model shows a statistically significant improvement over its base model. For each medical model in Table E2, we boldface the win rate if it wins more than it loses to its general-domain base model, and vice versa. As discussed in Section 3.1, we exclude the EHRNoteQA and i2b2 datasets from few-shot prompting evaluations, as most of the clinical notes already occupy the full context window for most models even in the zero-shot setting. Additionally, we exclude the zero-shot results on these datasets for MediTron-7B (and therefore its base model Llama-2-7B), as its context window size of 2k tokens is insufficient even in the zero-shot setting.
Overall, we observe that our findings discussed in Section 4.1.2 also hold in the constrained decoding setting. In particular, we observe that after performing model-specific prompt selection for each model independently,
@@
    only two medical LLMs—Med42-v1-70B and Med42-v2-8B—consistently show improvements in both zero-shot and 3-shot settings;
@@
E.3 Evaluation of Medical VLMs on Visual Medical QA (Section 4.1.3)
Figure E3: After independently selecting the best prompt format and examples for each model, medical VLMs do not consistently show a statistically significant improvement over their general-domain counterparts on visual QA tasks, in both (a) zero-shot and (b) 3-shot settings. The top row of each panel shows the absolute exact-match accuracies on the test set, and the bottom row of each panel shows the relative exact-match accuracies along with 95% confidence intervals derived via bootstrapping on the test set (see Section 3). Here, we show the results when model predictions are generated via constrained decoding. The results for greedy decoding are shown in Figure 5.
Table E3: The zero-shot and 3-shot win, tie, and loss rates (%) of all medical VLMs on visual QA, after independently optimizing the prompt for each model. For each medical model, we boldface the win rate if it wins more than it loses to its general-domain base model, and vice versa. Here, we show the results when model predictions are generated via constrained decoding. The results for greedy decoding are shown in Table 4.
@@
In Figure E3, we show the absolute and relative exact-match accuracies achieved by the medical and general-domain VLMs on the visual QA datasets, from zero-shot (Figure E3(a)) and 3-shot prompting (Figure E3(b)), respectively. In Table E3, we also show the win, tie, and loss rates (%) of the medical VLMs, where win rate refers to the proportion of visual QA datasets where a medical model shows a statistically significant improvement over its base model. For each medical model in Table E3, we boldface the win rate if it wins more than it loses to its general-domain base model, and vice versa.
Overall, we observe that our findings discussed in Section 4.1.3 also hold in the constrained decoding setting. In particular, we observe that after performing model-specific prompt selection for each model independently,
@@
    the absolute accuracies do not consistently exhibit an increasing trend going from the zero-shot setting to the 3-shot setting;
@@
E.4 Ignoring Prompt Sensitivity and Statistical Uncertainty Can Overestimate the Performance Benefits from Medical DAPT (Section 4.2.3)
@@
Figure E4: Optimizing the prompt for only the medical model and comparing models without accounting for statistical uncertainty can overestimate the performance improvements from medical DAPT. We show the win/tie/loss rates (%) of the medical models across all (model pair, QA dataset) combinations, when (a) independently optimizing the prompt for each model and performing statistical testing, (b) optimizing the prompt only for the medical model and performing statistical testing, (c) independently optimizing the prompt for each model without statistical testing, and (d) optimizing the prompt only for the medical model without statistical testing. Here, we show the results for constrained decoding. The results for greedy decoding are shown in Figure 8.

### v2

@@ C.2 Default LLM Prompt Formats
Here, we share the default prompt formats for each LLM, with MMLU (Clinical Knowledge) (Hendrycks et al., 2021) as a running example. We denote the system prompt in orange, the few-shot examples in green, and the question in pink. For models without a specific system prompt and prompt format designed for closed-ended medical QA (Section C.1), we use a manually designed prompt format by default. This includes all general-domain LLMs. For example, the default 1-shot prompt for non-instruction-tuned models is as follows:
@@
For instruction-tuned models, which expect a specific conversational format, we apply the above format to each “user” query and “assistant” response and remove the ‘###’ and ‘Answer:’ tags. For example, the input prompt to Llama-3-70B-Instruct is as follows:
@@
<|begin_of_text|><|start_header_id|> system <|end_header_id|>
@@
<|system|>: You are a helpful medical assistant created by M42 Health in the UAE. <|prompter|>: Question: Glycolysis is the name given to the pathway involving the conversion of:
@@ C.3 Default VLM Prompt Formats
Here, we share the default prompt formats that we use for each VLM, using VQA-RAD (Lau et al., 2018) as a running example. We denote the system prompt in orange, the few-shot examples in green, and the question pink. We show the format in the 1-shot setting.
@@
Appendix D Zero-/Few-Shot Prompting Evaluation Results with Constrained Decoding
@@
Table D1: The zero-shot and 3-shot win/tie/loss rates (%) of all medical LLMs on textual medical knowledge QA, after independently optimizing the prompt for each model. For each medical model, we boldface the win rate if it wins more than it loses to its general-domain base model, and vice versa. Here, we show the results when model predictions are generated via constrained decoding. The results for greedy decoding are shown in Table 3.
@@
OpenBioLLM-70B 7.7 76.9 15.4 23.1 53.8 23.1
MediTron-70B 30.8 46.2 23.1 15.4 69.2 15.4
Clinical-Camel-70B 18.2 72.7 9.1 9.1 72.7 18.2
OpenBioLLM-8B 0 53.8 46.2 7.7 69.2 23.1
MediTron-7B 23.1 76.9 0 7.7 92.3 0
BioMistral-7B 30.8 69.2 0 7.7 92.3 0
BioMedGPT-LM-7B 7.7 84.6 7.7 7.7 69.2 23.1
Aggregate 16.9 68.5 14.6 11.2 74.2 14.6
Here, we provide the constrained decoding results for the zero-/few-shot prompting evaluations described in Section 3.1 (see “Models” in Section 3 for a description on constrained decoding). In Appendices D.1D.3, corresponding to Sections 4.1.14.1.3, we show the results on the textual medical knowledge QA, textual clinical note QA, and visual medical QA datasets, when model-specific prompt selection is performed for each model independently. In Appendix D.4, corresponding to Section 4.2.3, we show the aggregate results on how the win, tie, and loss rates (%) of the medical models are affected by the inclusion (or lack thereof) of model-specific prompt selection and statistical testing.
D.1 Evaluation of Medical LLMs on Textual Medical Knowledge QA (Section 4.1.1)
Figure D1: After independently selecting the best prompt format and examples for each model, medical LLMs (textured bars) fail to consistently improve over their base models (solid bars) on textual medical knowledge QA tasks, in both (a) zero-shot and (b) 3-shot settings. In each panel, the top row shows the absolute exact-match accuracies on the test set, and the bottom row shows the relative exact-match accuracies along with 95% confidence intervals derived via bootstrapping on the test set (Section 3). Here, model predictions are generated via constrained decoding. Greedy decoding results are shown in Figure 4.
In Figures D1(a) and (b), we show the absolute and relative exact-match accuracies achieved by the medical and general-domain LLMs on the textual medical knowledge QA datasets, from zero-shot and 3-shot prompting, respectively. In Table D1, we also show the win, tie, and loss rates (%) of the medical LLMs, where win rate refers to the proportion of QA datasets where a medical model shows a statistically significant improvement over its base model. We exclude the results for Clinical-Camel-70B on both versions of MedQA, and the Med42 models on all textual medical knowledge QA datasets, as they have already been trained on these datasets via medical DAPT (Table 2).
Overall, we observe that our findings discussed in Section 4.1.1 also hold in the constrained decoding setting. In particular, we observe that
@@
    more recent general-domain models (e.g., Llama-3-8B) tend to show stronger performance even without any medical adaptation via DAPT; and
@@
    the majority of medical LLMs show limited improvement over their base models, reaching a tie in most cases (aggregate zero-shot/3-shot tie rates of 68.5%/74.2%).
@@
D.2 Evaluation of Medical LLMs on Textual Clinical Note QA (Section 4.1.2)
Figure D2: After independently selecting the best prompt format and examples for each model, medical LLMs (textured bars) fail to consistently improve over their base models (solid bars) on textual clinical note QA tasks, in both (a) zero-shot and (b) 3-shot settings. In the 3-shot setting, we exclude the results on the EHRNoteQA and i2b2 datasets given the context window limitations (Section 3.1). In each panel, the top row shows the absolute exact-match accuracies on the test set, and the bottom row shows the relative exact-match accuracies along with 95% confidence intervals derived via bootstrapping on the test set (Section 3). Here, model predictions are generated via constrained decoding. Greedy decoding results are shown in Figure 5.
Figure D3: Even after accounting for the imbalance in the distribution of clinical acronyms in the CASI and MIMIC-III datasets, medical LLMs (textured bars) fail to consistently improve over their base models (solid bars), in both (a) zero-shot and (b) 3-shot settings. We show the results when the prompts are optimized for each model independently. In each panel, the top row shows the absolute macro exact-match accuracies and F1 scores—averaged over clinical acronyms—on the test set, and the bottom row shows the relative macro exact-match accuracies and F1 scores along with 95% confidence intervals derived via bootstrapping on the test set (Section 3). Here, model predictions are generated via constrained decoding. Greedy decoding results are shown in Figure 6.
Table D2: The zero-shot and 3-shot win, tie, and loss rates (%) of all medical LLMs on textual clinical note QA, after independently optimizing the prompt for each model. For each medical model, we boldface the win rate if it wins more than it loses to its general-domain base model, and vice versa. Here, we show the results when model predictions are generated via constrained decoding. The results for greedy decoding are shown in Table 4.
@@
Med42-v2-70B 12.5 75.0 12.5 0 66.7 33.3
OpenBioLLM-70B 0 75.0 25.0 0 33.3 66.7
Med42-v1-70B 87.5 0 0 66.7 33.3 0
MediTron-70B 12.5 75.0 12.5 0 33.3 66.7
Clinical-Camel-70B 87.5 12.5 0 33.3 0 66.7
Med42-v2-8B 37.5 37.5 25.0 66.7 0 33.3
OpenBioLLM-8B 37.5 25.0 37.5 66.7 0 33.3
MediTron-7B 0 66.7 33.3 33.3 66.7 0
BioMistral-7B 0 12.5 87.5 33.3 0 66.7
BioMedGPT-LM-7B 50.0 12.5 37.5 33.3 0 66.7
Aggregate 34.7 37.3 28.0 33.3 23.3 43.3
In Figures D2(a) and (b), we show the absolute and relative exact-match accuracies achieved by the medical and general-domain LLMs on the textual clinical note QA datasets, from zero-shot and 3-shot prompting, respectively. In Figures D3(a) and (b), we show the absolute and relative macro exact-match accuracies and F1 scores—averaged over clinical acronyms—on the CASI and MIMIC-III sense disambiguation datasets, from zero-shot and 3-shot prompting, respectively. In Table D2, we also show the win, tie, and loss rates (%) of the medical LLMs, where win rate refers to the proportion of QA datasets where a medical model shows a statistically significant improvement over its base model. As discussed in Section 3.1, we exclude the EHRNoteQA and i2b2 datasets from few-shot prompting evaluations, as most of the clinical notes already occupy the full context window for most models even in the zero-shot setting. Additionally, we exclude the zero-shot results on these datasets for MediTron-7B (and its base model Llama-2-7B), as its context window size of 2k tokens is insufficient even in the zero-shot setting.
Overall, we observe that our findings discussed in Section 4.1.2 also hold in the constrained decoding setting. In particular, we observe that
@@
    only two medical LLMs—Med42-v1-70B and Med42-v2-8B—consistently show improvements in both zero-shot and 3-shot settings; and
@@
D.3 Evaluation of Medical VLMs on Visual Medical QA (Section 4.1.3)
Figure D4: After independently selecting the best prompt format and examples for each model, medical VLMs (textured bars) fail to consistently improve over their base models (solid bars) on visual medical QA tasks, in both (a) zero-shot and (b) 3-shot settings. In each panel, the top row shows the absolute exact-match accuracies on the test set, and the bottom row shows the relative exact-match accuracies along with 95% confidence intervals derived via bootstrapping on the test set (Section 3). Here, model predictions are generated via constrained decoding. Greedy decoding results are shown in Figure 7.
Table D3: The zero-shot and 3-shot win, tie, and loss rates (%) of all medical VLMs on visual medical QA, after independently optimizing the prompt for each model. For each medical model, we boldface the win rate if it wins more than it loses to its general-domain base model, and vice versa. Here, we show the results when model predictions are generated via constrained decoding. The results for greedy decoding are shown in Table 5.
@@
In Figures D4(a) and (b), we show the absolute and relative exact-match accuracies achieved by the medical and general-domain VLMs on the visual medical QA datasets, from zero-shot and 3-shot prompting, respectively. Table D3 shows the win, tie, and loss rates (%) of the medical VLMs, where win rate refers to the proportion of visual medical QA datasets where a medical model shows a statistically significant improvement over its base model.
Overall, we observe that our findings discussed in Section 4.1.3 also hold in the constrained decoding setting. In particular, we observe that
@@
    the absolute accuracies do not consistently exhibit an increasing trend going from the zero-shot setting to the 3-shot setting; and
@@
D.4 Ignoring Prompt Sensitivity and Statistical Uncertainty Can Overestimate the Performance Benefits from Medical DAPT (Section 4.2.3)
@@
Figure D5: Optimizing the prompt for only the medical model and comparing models without accounting for statistical uncertainty can overestimate the performance improvements from medical DAPT. We show the win/tie/loss rates (%) of the medical models across all (model pair, QA dataset) combinations, when (a) optimizing the prompt for each model, with statistical testing; (b) optimizing the prompt only for the medical model, with statistical testing; (c) optimizing the prompt for each model, without statistical testing; and (d) optimizing the prompt only for the medical model, without statistical testing. Here, we show the results for constrained decoding. Greedy decoding results are shown in Figure 10.